OpenStudy (anonymous):

Find the value of three times the quantity x minus two when x equals eight. 1 eighteen twenty-two twenty-four

OpenStudy (midhun.madhu1987):

3x - 2 x=8

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ 3(x-2) }\) When \(\normalsize\color{black}{ x=8}\) \(\normalsize\color{black}{ 3(x-2) }\) \(\normalsize\color{black}{ 3(8-2) }\) \(\normalsize\color{black}{ 3(6) }\) \(\normalsize\color{black}{ 18 }\)

OpenStudy (anonymous):

what

OpenStudy (anonymous):

im 13 i dont get it

OpenStudy (mrnood):

@SolomonZelman I would have read this as 3x - 2 (=22) (since multiply normally precedes addition)

OpenStudy (solomonzelman):

It says `times the QUANTITY of x-2`

OpenStudy (anonymous):

what is the anwser

OpenStudy (anonymous):

can u fan me

OpenStudy (anonymous):

3(x)-2=? 3(8)-2=? 24-2= 22 22 is your answer

OpenStudy (anonymous):

thank u

OpenStudy (solomonzelman):

@Miahkw, it is \(\normalsize\color{black}{3(x-2) }\) and NOT \(\normalsize\color{red}{3(x)-2 }\)

OpenStudy (mrnood):

I think this is open to debate, unfortunately BOTH possible answers are in the list.

OpenStudy (mrnood):

@SolomonZelman NOt trying to start a flame! It COULD be read as "three times the quantity x" - 2